Solve each of the following equations:

(i) e2x Β· (e2x-1)3 = 9

(ii) 2 ln x - ln(3 - 5x) = ln 2


1
Expert's answer
2021-09-26T18:04:27-0400

(i)


"e^{2x}\cdot(e^{2x}-1)^3=9"

Use substitution "e^{2x}=t, t>0." Then


"t(t-1)^3=9"

"t^4-3t^3+3t^2-t-9=0"



"t>0=>t\approx2.527"


"e^{2x}\approx2.527"

"x\approx0.5\ln2.527"

"x\approx0.464"



(ii)


"2\ln x-\ln (3-5x)=\ln 2"

"x>0, 3-5x>0"

Then "0<x<\dfrac{3}{5}"


"\ln x^2=\ln(2(3-5x))"

"x^2=6-10x"

"x^2+10x-6=0"

"D=(10)^2-4(1)(-6)=124"

"x=\dfrac{-10\pm\sqrt{124}}{2(1)}=-5\pm\sqrt{31}"

"-5+\sqrt{31}<\dfrac{3}{5}"

"\sqrt{31}<\dfrac{28}{5}"

"31<\dfrac{784}{25}"

"31<31.36, True"

Answer: "x=\sqrt{31}-5."